### v1

@@ 3 Methodology
As illustrated in Fig.3, FOCUS is a unified large vision-language model that integrates pixel-level perception and controllable image generation in an end-to-end framework. The model comprises four core components, covering the entire process from visual encoding to image synthesis.  Section 3.1 introduce a dual-branch visual encoder and a generation-oriented visual tokenizer are employed to extract global semantic features and multi-scale fine-grained representations from inputs at different resolutions, discretizing the visual content into high-level tokens suitable for downstream editing tasks. Next, in the  Section 3.2 a large language model based on Qwen2.5 is introduced to unify multimodal input formats through task prompts, enabling support for a wide range of vision-language tasks. Building on this, in the  Section 3.3 a segmentation decoder is designed to perform high-precision fine-grained object segmentation. A diffusion-based image generator then synthesizes high-fidelity images guided by spatial segmentation masks and conditioned on discrete visual tokens in the  Section 3.4. In the  Section 3.5 further details the progressive training strategy and the construction of multi-source datasets.
@@ 3.1 Dual Visual Encoders and Generative Tokenizer
Specifically, we use the fused dual-branch features $E^{\prime}_{\text{img}}$ as the global semantic representation for generation, which are quantized into discrete semantic tokens and decoded through a lightweight decoder. During the visual tokenizer’s pretraining stage, this semantic branch is supervised with a cosine similarity loss against the original encoder features. A downsampling rate of 28×28$\times$28 × is adopted to align with mainstream vision-language models and focus on high-level semantic concepts.
We further extend the generative visual tokenizer with a pixel branch, following the standard MoVQGAN design. This branch uses a 16×16$\times$16 × downsampling rate to preserve fine textures. After quantization, the semantic and pixel tokens are concatenated along the channel dimension and passed to the decoder for image reconstruction. The pixel branch is trained with a combination of L1 loss, perceptual loss, and adversarial loss. We use large codebooks for both branches, with a size of 32,768 for the semantic branch and 98,304 for the pixel branch.
@@ 4.1 Implementation Details
Method LLM POPE MMBench SEED MME-P MM-Vet MMMU AI2D VQA-text ChartQA DocVQA InfoVQA OCRBench
@@ 4.3 Controllable Generation and Editing
FOCUS AR Und, Gen, Edit 0.826 0.872 0.275 0.101
@@ Appendix B Progressive Large Vision Language Model Training
Visual-Language Projector Warmup. This stage establishes initial alignment between visual features and the large language model (LLM). We train only the projection heads of the vanilla and pixel encoders, along with the learnable queries and gated cross-attention modules in the hierarchical encoder. All backbone encoders, the LLM, the mask decoder is Mask2Former [15], and the diffusion decoder are frozen. Training is performed on 256$\times$256 image-text pairs using a standard language modeling objective, without any segmentation or generation supervision.
Multimodal Pretraining with Segmentation-Aware Alignment. We activate and jointly train the LLM, visual-language adapters, and the mask decoder to enhance multimodal understanding and segmentation capability. The diffusion decoder remains frozen. Supervision includes token-level cross-entropy, segmentation losses (Dice and BCE), and visual reconstruction losses using precomputed features. Training is conducted at both 256$\times$256 and 512$\times$512 resolutions on datasets covering image-text alignment, referring segmentation, and multimodal reasoning.
Instruction Tuning for Region-Controlled Editing. This stage focuses on segmentation perception and controllable generation for LVLMs. We instruct tuning the LLM, the mask decoder, and the cross-attention layers in the diffusion decoder, while freezing all visual encoders. Images are processed at resolutions up to 1024$\times$1024. Segmentation outputs are used as spatial guidance and injected into the diffusion decoder’s attention layers. Supervision includes text generation loss, segmentation loss, and L2 loss between denoised outputs and targets.
@@
Stage I: Visual Quantizer and Diffusion Pretraining 45M Image-to-Image COYO [6], EMOVA [9], LAION-2B [64]
Stage II: Visual-Language Projector Warmup 30M Image-to-Text LLaVA-150K [38, 37], COYO, EMOVA-Pretrain
Stage III: Multimodal Pretraining with Generative and Structural Signals 35M Image-to-Text & Editing UltraEdit [84], AnyEdit [78], SEED-Edit [55]
@@
Stage IV: Instruction Tuning for Controllable Editing 7M Image Editing EMOVA-SFT, Pixmo [18], M4-Instruct [37],
@@ Appendix D Training Configurations Across Different Stages
Learning Rate 1e-4 (semantic) 2e-4 (pixel) 2e-5 1e-3 2e-5 (Visual encoder, LLM) 1e-3(Mask Decoder) 2e-5 (Visual encoder, LLM) 2e-6 (Mask Decoder, Diffusion)
@@
Training Steps 136k (pixel) 28k (semantic) 220k 1epoch 3epoch 1epoch
@@
Frozen Modules Vanilla encoder Hierarchical Encoder All encoders Codebooks Visual Encoder, LLM Mask Decoder, Diffusion Diffusion Visual Encoder

### v2

@@ 3 Methodology
As illustrated in Fig.3, FOCUS is a unified large vision-language model that integrates pixel-level perception and controllable image generation in an end-to-end framework. The model comprises four core components, covering the entire process from visual encoding to image synthesis.  Section˜3.1 introduce a dual-branch visual encoder and a generation-oriented visual tokenizer are employed to extract global semantic features and multi-scale fine-grained representations from inputs at different resolutions, discretizing the visual content into high-level tokens suitable for downstream editing tasks. Next, in the  Section˜3.2 a large language model based on Qwen2.5 is introduced to unify multimodal input formats through task prompts, enabling support for a wide range of vision-language tasks. Building on this, in the  Section˜3.3 a segmentation decoder is designed to perform high-precision fine-grained object segmentation. A diffusion-based image generator then synthesizes high-fidelity images guided by spatial segmentation masks and conditioned on discrete visual tokens in the  Section˜3.4. In the  Section˜3.5 further details the progressive training strategy and the construction of multi-source datasets.
@@ 3.1 Dual Visual Encoders and Generative Tokenizer
Specifically, we use the fused dual-branch features $E^{\prime}_{\text{img}}$ as the global semantic representation for generation, which are quantized into discrete semantic tokens and decoded through a lightweight decoder. During the visual tokenizer’s pretraining stage, this semantic branch is supervised with a cosine similarity loss against the original encoder features. A downsampling rate of $28\times$ is adopted to align with mainstream vision-language models and focus on high-level semantic concepts.
We further extend the generative visual tokenizer with a pixel branch, following the standard MoVQGAN design. This branch uses a $16\times$ downsampling rate to preserve fine textures. After quantization, the semantic and pixel tokens are concatenated along the channel dimension and passed to the decoder for image reconstruction. The pixel branch is trained with a combination of L1 loss, perceptual loss, and adversarial loss. We use large codebooks for both branches, with a size of 32,768 for the semantic branch and 98,304 for the pixel branch.
@@ 4.1 Implementation Details
Method LLM POPE MMBench SEED MME-P $M$-Vet MMMU AI2D VQA-text ChartQA DocVQA InfoVQA OCRBench
@@ 4.3 Controllable Generation and Editing
ILLUME+ AR Und, Gen, Edit 0.826 0.872 0.275 0.101
FOCUS AR Und, Gen, Edit 0.831 0.876 0.278 0.105
@@ Appendix B Progressive Large Vision Language Model Training
Visual-Language Projector Warmup. This stage establishes initial alignment between visual features and the large language model (LLM). We train only the projection heads of the vanilla and pixel encoders, along with the learnable queries and gated cross-attention modules in the hierarchical encoder. All backbone encoders, the LLM, the mask decoder is Mask2Former [15], and the diffusion decoder are frozen. Training is performed on $\times256$ image-text pairs using a standard language modeling objective, without any segmentation or generation supervision.
Multimodal Pretraining with Segmentation-Aware Alignment. We activate and jointly train the LLM, visual-language adapters, and the mask decoder to enhance multimodal understanding and segmentation capability. The diffusion decoder remains frozen. Supervision includes token-level cross-entropy, segmentation losses (Dice and BCE), and visual reconstruction losses using precomputed features. Training is conducted at both $\times256$ and $\times512$ resolutions on datasets covering image-text alignment, referring segmentation, and multimodal reasoning.
Instruction Tuning for Region-Controlled Editing. This stage focuses on segmentation perception and controllable generation for LVLMs. We instruct tuning the LLM, the mask decoder, and the cross-attention layers in the diffusion decoder, while freezing all visual encoders. Images are processed at resolutions up to $\times1024$. Segmentation outputs are used as spatial guidance and injected into the diffusion decoder’s attention layers. Supervision includes text generation loss, segmentation loss, and L2 loss between denoised outputs and targets.
@@
Stage I: Visual Quantizer and
Diffusion Pretraining
45M Image-to-Image COYO [6], EMOVA [9], LAION-2B [64]
Stage II: Visual-Language
Projector Warmup
30M Image-to-Text LLaVA-150K [38, 37], COYO, EMOVA-Pretrain
Stage III: Multimodal Pretraining
with Generative and Structural Signals
35M Image-to-Text & Editing UltraEdit [84], AnyEdit [78], SEED-Edit [55]
@@
Stage IV: Instruction Tuning
for Controllable Editing
7M Image Editing EMOVA-SFT, Pixmo [18], M4-Instruct [37],
@@ Appendix D Training Configurations Across Different Stages
Learning Rate
1e-4 (semantic)
2e-4 (pixel)
2e-5
1e-3
2e-5 (Visual encoder, LLM)
1e-3(Mask Decoder)
2e-5 (Visual encoder, LLM)
2e-6 (Mask Decoder, Diffusion)
@@
Training Steps
136k (pixel)
28k (semantic)
220k 1epoch 3epoch 1epoch
@@
Frozen Modules
Vanilla encoder
Hierarchical Encoder
All encoders
Codebooks
Visual Encoder, LLM
Mask Decoder, Diffusion
Diffusion
Visual Encoder